A triangle has side lengths of 21,29 and 20 what is the area of incenter circle ???
Answer by richard1234(7193)

Triangle is a right triangle since 20^2 + 21^2 = 29^2. Its area is (1/2)(20)(21) = 210.
Using area A = rs (inradius r, semiperimeter s), we have 210 = 35r, or r = 6. Then the area of the inscribed circle is 36pi.